OpenStudy (thivitaa):

Parametic equation of a curve are given X=at² and y=at³ where a is a constant. Find Equation of tangent at point where t=2. Hence find coordinates where tangent meets curve again.

OpenStudy (anonymous):

\[x^3=a^3t^6 \implies y^2=a^2t^6\] \[x^3=ay^2\] \[2a y y'=3x^2\] when \[t=2,x=2^6a^3,y^2=a^22^6\] find y'

OpenStudy (anonymous):

\[y'=\frac{dy}{dx}=\frac{\frac{dy}{dt}}{\frac{dx}{dt}}=\frac{2at}{3at^2}=\frac{2}{3}t~\text{ for }~t\not=0\] Slope of tangent line at \(t=2\): \[y'(2)=\frac{2}{3}\times 2=\frac{4}{3}\] Point on curve when \(t=2\): \[\begin{cases}x(2)=a\times2^2\\y(2)=a\times2^3\end{cases}~\Rightarrow~(x,y)=(4a,8a)\] So now find the equation of the tangent line that passes through \((4a,8a)\) with slope \(\dfrac{4}{3}\), then find the intersection with the curve for \(t\not=2\).